# Foliations by Critical Surfaces of the Hawking Energy in Asymptotically Flat Initial Data Sets

### v1

@@ 2.2. Large-sphere limit
Before continuing, we will need the following result from [24, Proposition A1]:
@@
$|z|\leq cR,\qquad\max_{\Sigma}\left|\nu-\frac{x-z}{R}\right|\leq cR^{\frac{1+\epsilon}{2}},\qquad|\Sigma|\leq cR^{2},$
@@
$\left|\,E_{ADM}+\frac{R}{8\pi}\int_{\Sigma}\bigl{(}\mathrm{Ric}(\nu,\nu)-\frac{\mathrm{Sc}}{2}\bigr{)}\,d\mu\right|\leq\frac{C}{R^{\epsilon}},$
where $R.\!=\min_{\Sigma}|x|$, and $\nu$, $\mu$ denote the outer unit normal and the surface measure of $\Sigma\to M$, respectively.
For more information on this method to calculate the ADM energy see [13, 23]. This applies to both constructions: it holds for the leaves of our foliation and for the family of Theorem 2.14 (with the choice $z=r\xi$). We now prove:
@@
By using Proposition 2.17 and the Gauss equation we have.
$\begin{split}|m-\mathcal{E}(\Sigma)|\leq&\left|m-\frac{R}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu\right|+\left|\frac{R}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu-\mathcal{E}(\Sigma(r))\right|\\ \leq&\left|\frac{R}{16\pi}\int_{\Sigma(r)}\left(\mathrm{Sc}^{\Sigma}-\frac{H^{2}}{2}+|\mathring{B}|^{2}\right)d\mu-\sqrt{\frac{|\Sigma(r)|}{16\pi}}\left(1-\frac{1}{16\pi}\int_{\Sigma(r)}(H^{2}-P^{2})\,d\mu\right)\right|\\ &+\mathcal{O}(r^{-\epsilon})\end{split}$
@@
for some $\epsilon>0$, and where $\mathrm{Sc}^{\Sigma}$ is the scalar curvature of $\Sigma$. Now by Gauss-Bonnet theorem, the decay on $k$ and that $r=R+\mathcal{O}(r^{-1})$ we obtain
@@
(48) $|m-\mathcal{E}(\Sigma(r))|\leq\frac{r}{16\pi}\int_{\Sigma(r)}|\mathring{B}|^{2}d\mu+\mathcal{O}(r^{-\epsilon}).$
Note that our spheres satisfy $H=\frac{2}{r}+\mathcal{O}(r^{-\epsilon})$ and therefore we can directly apply the bootstrap argument of [4, Proposition 1] or [25, Proposition 3.7] and obtain that $\mathring{B}||_{L^{2}(\Sigma(r))}=\mathcal{O}(r^{-\frac{1}{2}-\epsilon})$ for some $\epsilon>0$, then the result follows. ∎
@@
With this result, we have established that the Hawking energy evaluated on Hawking surfaces satisfies the large-sphere limit. A study of the small-sphere limit on such surfaces can be found in [11, 26].
@@
Note that this large-sphere limit converges to the ADM energy and not the ADM mass $\sqrt{E_{ADM}^{2}-p_{ADM}^{2}}$, the reason for this is that when considering asymptotically flat manifolds (under standard decay assumptions), the decay of $k$ is too rapid to capture the momentum contribution in the limit. This not only happens with this foliation, the same occurs for the STCMC foliation of Cederbaum and Sakovich, see [4, Proposition 5.6]. For this reason, we consistently view $\mathcal{E}(\Sigma)$ here as a quasi-local energy, not a mass.
@@ 3. Center of the foliation
We now study the center of the foliations constructed above. Because the foliation is composed of on-center spheres, any natural geometric center lies in the compact interior region $K$ from Definition 1.2, which is not visible through the asymptotic coordinate chart. For this reason, we introduce instead the coordinate center of the foliation, defined at infinity via the ambient chart.
Definition 3.1 (Coordinate center of a foliation).
$\vec{C}:=\lim_{\sigma\to\infty}\vec{c}(\Sigma_{\sigma}),$
The study of the center of geometric foliations in asymptotically flat manifolds has attracted substantial attention. The first major result in this direction was due to Huisken and Yau [17]. In an asymptotically flat, time-symmetric initial data set they constructed a unique foliation of the end by constant mean curvature (CMC) spheres. Under additional asymptotic symmetry assumptions, the coordinate center of this foliation coincides with the Hamiltonian center of mass $\vec{C}_{\mathrm{H}}=(C^{1}_{\mathrm{H}},C^{2}_{\mathrm{H}},C^{3}_{\mathrm{H}})$, which is given by
@@
where $E$ is the ADM energy and $d\mu_{\delta}$ is the Euclidean volume element on the coordinate sphere of radius $r$. See [29, 1] for further discussion of this formulation and its invariance properties.
This construction was generalized to the dynamical setting by Cederbaum and Sakovich in [4], who constructed a unique foliation by STCMC (spacetime constant mean curvature) surfaces in appropriate asymptotically flat initial data sets. They proved that the coordinate center of that foliation converges, and identified its limit as a suitably generalized center of mass—the STCMC center of mass—which reduces to the Huisken–Yau/Hamiltonian center in the time-symmetric case.
Eichmair and Koerber in [8] showed that the center of the Willmore foliation captures the Hamiltonian center of mass.
Theorem 3.2 ([8, Theorem 2]).
@@
It is then natural to ask whether the center of the foliation by Hawking surfaces similarly encodes the appropriate center of mass in the dynamical setting—in particular, the STCMC center of mass constructed in [4]. To address this, we study the coordinate center of the foliation produced in Theorem 2.10.
As observed earlier, such foliations need not be centered in general. This indicates that centering requires stronger asymptotic symmetry and decay. Accordingly, we impose the following additional hypotheses on the scalar curvature and the tensor $k$:
@@
Here, for any continuous function $f$ on $\mathbb{R}^{3}$, we denote its even and odd parts by
and for tensors we take even/odd parts componentwise in the chosen chart. These symmetry assumptions are a direct consequence of the Regge-Teitelboim conditions introduced in [29]. In particular, they guarantee the convergence of the Hamiltonian center of mass.
Following Cederbaum–Sakovich [4], the spacetime (STCMC) center of mass $C_{\mathrm{STCMC}}$ extends the Hamiltonian notion to the dynamical case. It can be written as the Hamiltonian center of mass plus the $k$-dependent correction
(54) $\lim_{r\to\infty}\frac{r^{2}}{32\pi E_{ADM}}\int_{S_{r}(0)}\big{(}\pi(\nu,\nu)\big{)}^{2}\,\nu\,d\mu_{\delta},$
Under (52) this correction term vanishes.
@@
using $x^{i}\partial_{i}(|x|^{2}R)\leq\mathcal{O}(|x|^{-3})$, the parity conditions and using the same argument as in [8, Lemma 11] (or [9, Lemma 24]) we have
@@
Note that without this result the foliation cannot be centered.
We have seen that each leaf $\Sigma_{\xi,r}$ of the foliation can be expressed as the graph over $S_{\xi,r}$ of a function $u=-2+\mathcal{O}(|\xi|^{2})$, this implies that when comparing $\Sigma_{\xi,r}$ with $S_{\xi,r}$ the difference would be given by $2+\mathcal{O}(|\xi|^{2})$. The comparison between the two surfaces will be important when trying to study the center of the surfaces $\Sigma_{\xi,r}$ in terms of the center of the spheres $S_{\xi,r}$, that is why it is more practical to consider the spheres $S_{\xi,r}$ with a different radius $\tilde{r}=r-2$ and have the difference between $\Sigma_{\xi,r}$ and $S_{\xi,r}$ be given by the error term $\mathcal{O}(|\xi|^{2})$. Summarizing, following [8] we define:
@@
With these definitions and the previous results, we obtain the following result, which is the same as [8, Lemma 13] and has the same proof.
@@
We also have the following result equivalent to [8, Lemma 14]
@@
Note that (60) is the same expression obtained in [8, Lemma 14] and (61) is obtained directly by using the decay of $k$. This shows that we can reuse many of the results of [8] concerning the expression (60). We can then apply directly [8, Proposition 18] to obtain
@@
After a lengthy calculation it was shown in [8, Proposition 18] that the second term of the previous expression is related to the Hamiltonian center of mass. Then replacing the second term and the variation $\bar{D}\int_{S_{\tilde{\xi},\tilde{r}}}P^{2}\,d\mu_{\delta}$, we arrive at the equivalent of [8, Corollary 19].
@@
Let $(M,g,k)$ be $\mathcal{C}^{4}$-asymptotic to Schwarzschild with mass $m>0$ satisfying the condition of Theorem 2.10. Furthermore, assume that
@@
Under the stronger decay and parity hypotheses (and in the time-symmetric case), Theorem 3.2 shows that the center of the Willmore foliation agrees with the Hamiltonian center of mass. In general, the Willmore center is sensitive to the asymptotic distribution of the scalar curvature and may differ from the Hamiltonian center unless additional strong asymptotic symmetry is imposed; this is evident from (63) with $k=0$ and was also observed by Eichmair–Koerber [8].
@@
$\begin{split}\frac{r^{4}}{128\pi}\int_{S_{r}(r\xi(r))}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}&=\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2DW_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}\\ &+\frac{r^{4}}{256\pi}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\delta(e_{i},\nu)d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi-q,r\xi-q)\end{split}$
Where $\tilde{\xi}$ is some intermediate point on the straight-line segment from $|q|$ to $\xi r$. Now as in Lemma 3.3 we see that the third term decays as $\mathcal{O}(|\xi|)=\mathcal{O}(r^{-1})$ obtaining
$\begin{split}\frac{r^{4}}{128\pi}\int_{S_{r}(r\xi(r))}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}=&\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2DW_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\end{split}$
@@
(69) $\begin{split}r\delta(\xi(r&),e_{i})=\delta(C_{H},e_{i})+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\Big{(}\delta(e_{i},\nu)\delta(r\xi-q,\nu)\bar{D}_{\nu}\mathrm{Sc}+3\delta(e_{i},\nu)\delta(r\xi-q,\nu)\frac{\mathrm{Sc}}{r}-\frac{\mathrm{Sc}}{r}\,\delta(e_{i},r\xi-q)\Big{)}d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2DW_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\end{split}$

### v2

@@ 2.2. Large-sphere limit
Before continuing, we will need the following result from [Nerz2, Proposition A1]:
@@
$|z|\leq cr_{min},\qquad\max_{\Sigma}\left|\nu-\frac{x-z}{\,r_{min}}\right|\leq cr_{min}^{-\frac{1+\epsilon}{2}},\qquad|\Sigma|\leq cr_{min}^{2},$
@@
$\left|\,E_{ADM}+\frac{r_{min}}{8\pi}\int_{\Sigma}\bigl(\mathrm{Ric}(\nu,\nu)-\frac{\mathrm{Sc}}{2}\bigr)\,d\mu\right|\leq\frac{C}{r_{min}^{\epsilon}},$
where $r_{min}:=\min_{\Sigma}|x|$, and $\nu$, $\mu$ denote the outer unit normal and the surface measure of $\Sigma\to M$, respectively.
The Ricci–flux identity for the ADM mass used here goes back to Ashtekar–Hansen [ashtekar1978unified] and Chruściel [chrusciel1986remark]; see also the modern formulations by Miao–Tam [miao2016evaluation] and Herzlich [herzlich2016computing].
We now prove:
@@
We begin by observing that the estimate from Proposition 2.17 may be applied in a translated asymptotically flat chart centered at the vector $z=r\xi(r)$ associated with the surface $\Sigma(r)$. Indeed, define the translated coordinates
$x^{\prime}:=x-z,\qquad\hat{r}:=\min_{x\in\Sigma}|x-z|.$
Since the asymptotically flat metric satisfies $g_{ij}(x)=\delta_{ij}+\mathcal{O}(|x|^{-1})$ and its derivatives decay as $\mathcal{O}(|x|^{-2})$, we have for large $|x^{\prime}|$ that $|x^{\prime}+z|\simeq|x^{\prime}|$ whenever $\frac{|z|}{|x^{\prime}|}\text{ remains bounded.}$
In our case $|z|=r|\xi|$ and $|\xi|\leq 1-\tilde{\delta}$, hence $|z|\leq(1-\tilde{\delta})r\leq C\,\hat{r}$, and therefore $\frac{|z|}{|x^{\prime}|}$ is uniformly bounded in the asymptotic region. Consequently, the metric remains asymptotically flat in the translated chart:
$g_{ij}(x^{\prime}+z)=\delta_{ij}+\mathcal{O}(|x^{\prime}|^{-1}),\qquad\partial g_{ij}(x^{\prime}+z)=\mathcal{O}(|x^{\prime}|^{-2}),$
and the same decay holds for higher derivatives. Thus, Proposition 2.17 applies with $\hat{r}$ in place of the original inner radius $r_{min}$. This translation merely recenters the coordinate system near the geometric center $z=r\xi$ of $\Sigma(r)$ and does not affect any asymptotic quantities such as the ADM energy.
We now check that the remaining hypotheses Proposition 2.17 are satisfied for the surface $\Sigma(r)$ in the translated chart, with $z=r\xi$ and $\hat{r}=\min_{\Sigma(r)}|x-z|$. First note that $\Sigma(r)$ is a normal graph of $u_{r}$ over $S_{r}(r\xi(r))$ and by Proposition 2.3, $\|u_{r}\|_{\mathcal{C}^{2}}$ is bounded independent of $r$ then
$\hat{r}=\min_{x\in\Sigma(r)}|x-r\xi(r)|\leq\|u_{r}\|_{\mathcal{C}^{0}}+r\leq Cr$
for some constant $C$. Given $x\in\Sigma(r)$ then there exit a $y\in S_{r}(r\xi(r))$ such that $x=y+u_{r}(y)\frac{y-r\xi(r)}{r}$ and
$r=|y-r\xi(r)|=|x-u_{r}(y)\frac{y-r\xi(r)}{r}-r\xi(r)|\leq|x-r\xi(r)|+\|u_{r}\|_{\mathcal{C}^{0}}\quad\forall x\in\Sigma(r).$
Then $r\leq\hat{r}+\|u_{r}\|_{\mathcal{C}^{0}}$ and there exist a constant $C^{\prime}$ such that $r\leq C^{\prime}\hat{r}$. This implies that $|r-\hat{r}|\leq\|u_{r}\|_{\mathcal{C}^{0}}$ and there is a constant $\tilde{C}$ such that $\frac{1}{\tilde{C}}\hat{r}\leq r\leq\tilde{C}\hat{r}$, therefore we can consider the two radius to be equivalent for our purposes.
Note that in these new coordinates the vector $z$ of the proposition is zero, then it follows directly that $|z|=0\leq\hat{r}$ and $|\Sigma(r)|\leq c\hat{r}^{2}$. It remains to check the condition on the normal vector. Note that since $\Sigma(r)$ is a graph over $S_{r}(r\xi(r))$, it holds that $\left|\nu-\frac{x-r\xi(r)}{r}\right|\leq Cr^{-1}\|u_{r}\|_{\mathcal{C}^{1}}$ then with this we can see that.
$\begin{split}\left|\nu-\frac{x^{\prime}}{\,\hat{r}}\right|&=\left|\nu-\frac{x-r\xi(r)}{\hat{r}}\right|\leq\left|\nu-\frac{x-r\xi(r)}{r}\right|+\left|\frac{x-r\xi(r)}{r}-\frac{x-r\xi(r)}{\hat{r}}\right|\\ &\leq Cr^{-1}\|u_{r}\|_{\mathcal{C}^{1}}+\left|\frac{1}{\hat{r}}-\frac{1}{r}\right||x-r\xi(r)|\leq Cr^{-1}\|u_{r}\|_{\mathcal{C}^{1}}+Cr^{-2}|x-r\xi(r)|\\ &\leq C\hat{r}^{-1}\end{split}$
Hence the assumptions of Proposition 2.17 hold for the translated surfaces $\Sigma(r)$, and we may apply its conclusion with $\hat{r}$ in place of $r_{min}$.
By Proposition 2.17 and that $|r-\hat{r}|\leq\|u_{r}\|_{\mathcal{C}^{0}}$ we have
(57) $\begin{split}m&=\frac{\hat{r}}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu+\mathcal{O}(r^{-\epsilon})\\ &=\frac{r}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu+\frac{\hat{r}-r}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu+\mathcal{O}(r^{-\epsilon})\\ &=\frac{r}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu+\mathcal{O}(r^{-\epsilon})\end{split}$
@@
Then by the Gauss equation, Gauss-Bonnet theorem and that by the decay of $k$ it holds $\sqrt{\frac{|\Sigma(r)|}{16\pi}}\int_{\Sigma(r)}P^{2}\,d\mu=\mathcal{O}(r^{-\epsilon})$, we have
@@
(58) $\begin{split}|m-\mathcal{E}(\Sigma)|\leq&\left|m-\frac{r}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu\right|+\left|\frac{r}{16\pi}\int_{\Sigma(r)}(\mathrm{Sc}-2\,\mathrm{Ric}(\nu,\nu))\,d\mu-\mathcal{E}(\Sigma(r))\right|\\ \leq&\left|\frac{r}{16\pi}\int_{\Sigma(r)}\left(\mathrm{Sc}^{\Sigma}-\frac{H^{2}}{2}+|\mathring{B}|^{2}\right)d\mu-\sqrt{\frac{|\Sigma(r)|}{16\pi}}\left(1-\frac{1}{16\pi}\int_{\Sigma(r)}H^{2}d\mu\right)\right|+\mathcal{O}(r^{-\epsilon})\\ \leq&\frac{r}{16\pi}\int_{\Sigma(r)}|\mathring{B}|^{2}d\mu+\mathcal{O}(r^{-\epsilon})\end{split}$
for some $\epsilon>0$.
To estimate $\int_{\Sigma(r)}|\mathring{B}|^{2}d\mu$ first note that since $\Sigma(r)$ is a Hawking surface it satisfies equation (4), which like in (13) and (14) can be written as $H\lambda+W_{1}+W_{2}=0$,and by the decay of $k$ it reduces to $H\lambda+W_{1}=\mathcal{O}(r^{-5})$. Dividing the equation by $H>0$, integrating over $\Sigma(r)$, and integrating by parts the term $\Delta_{\Sigma}H/H$ gives
(59) $\int_{\Sigma_{r}}\lambda+|\nabla^{\Sigma}\log H|^{2}+\frac{1}{2}|\mathring{B}|^{2}+\mathrm{Ric}(\nu,\nu)d\mu=\mathcal{O}(r^{-2}).$
Since $\lambda$ is nonnegative (see Lemma 2.5) and by the decay $\mathrm{Ric}=\mathcal{O}(r^{-3})$ we have
(60) $\int_{\Sigma(r)}|\mathring{B}|^{2}d\mu=\mathcal{O}(r^{-1})$
However this decay is not good enough to prove our result, in orther to improve the decacy we will use the the bootstrapping result of Cederbaum–Sakovich [STCMC, Proposition 4.5], which is an adaptation of the estimates in Nerz [Nerz, Proposition 2.4] to the spacetime mean–curvature (STCMC) setting. Roughly speaking, their result asserts that for a family of closed surfaces $\Sigma\subset M$ which are small normal graphs over coordinate spheres in an $\mathcal{C}^{2,\frac{1}{2}+\varepsilon}$-asymptotically flat manifold, if the mean curvature satisfies
$H=\frac{2}{r}+\mathcal{O}(r^{-1-\varepsilon}),$
and if $\|\mathring{B}\|_{L^{2}(\Sigma)}$ is sufficiently small, then one obtains the quantitative decay
$\|\mathring{B}\|_{L^{2}(\Sigma)}=\mathcal{O}(r^{-1/2-\varepsilon}),\qquad\|\mathring{B}\|_{L^{\infty}(\Sigma)}=\mathcal{O}(r^{-3/2-\varepsilon}),$
In our case, the spheres $\Sigma(r)$ constructed in Theorem 2.10 and Theorem 2.14 satisfy all these hypotheses: they are small normal graphs over $S_{r}(r\xi(r))$, the ambient metric $(M,g)$ is $\mathcal{C}^{2,1/2+\varepsilon}$–asymptotically flat, by the analysis above $H=\tfrac{2}{r}+\mathcal{O}(r^{-1-\varepsilon})$ and $\|\mathring{B}\|_{L^{2}(\Sigma(r))}=\mathcal{O}(r^{-\frac{1}{2}})$. Hence we can apply [STCMC, Proposition 4.5] directly to conclude that
$\|\mathring{B}\|_{L^{2}(\Sigma(r))}=\mathcal{O}(r^{-1/2-\varepsilon}).$
Substituting this decay into
$|m-\mathcal{E}(\Sigma(r))|\leq\frac{r}{16\pi}\!\int_{\Sigma(r)}|\mathring{B}|^{2}\,d\mu+\mathcal{O}(r^{-\epsilon})$
gives the desired convergence $\mathcal{E}(\Sigma_{r})\to m$ as $r\to\infty$. ∎
@@
With this result, we have established that the Hawking energy evaluated on Hawking surfaces satisfies the large-sphere limit. A study of the small-sphere limit on such surfaces can be found in [Alex, diaz2023local].
@@
Note that this large-sphere limit converges to the ADM energy and not the ADM mass $\sqrt{E_{ADM}^{2}-p_{ADM}^{2}}$, the reason for this is that when considering asymptotically flat manifolds (under standard decay assumptions), the decay of $k$ is too rapid to capture the momentum contribution in the limit. This not only happens with this foliation, the same occurs for the STCMC foliation of Cederbaum and Sakovich, see [STCMC, Proposition 5.6]. For this reason, we consistently view $\mathcal{E}(\Sigma)$ here as a quasi‐local energy, not a mass.
@@ 3. Center of the foliation
The study of the center of geometric foliations in asymptotically flat manifolds has attracted substantial attention. The first major result in this direction was due to Huisken and Yau [HY]. In an asymptotically flat, time-symmetric initial data set they constructed a unique foliation of the end by constant mean curvature (CMC) spheres. Under additional asymptotic symmetry assumptions, the coordinate center of this foliation coincides with the Hamiltonian center of mass $\vec{C}_{\mathrm{H}}=(C^{1}_{\mathrm{H}},C^{2}_{\mathrm{H}},C^{3}_{\mathrm{H}})$, which is given by
@@
where $E$ is the ADM energy and $d\mu_{\delta}$ is the Euclidean volume element on the coordinate sphere of radius $r$. See [beigmurch, regge1974role] for further discussion of this formulation and its invariance properties.
This construction was generalized to the dynamical setting by Cederbaum and Sakovich in [STCMC], who constructed a unique foliation by STCMC (spacetime constant mean curvature) surfaces in appropriate asymptotically flat initial data sets. They proved that the coordinate center of that foliation converges, and identified its limit as a suitably generalized center of mass—the STCMC center of mass—which reduces to the Huisken–Yau/Hamiltonian center in the time-symmetric case. The center of the STCMC foliation can be written as the Hamiltonian center of mass plus the $k$-dependent correction
(62) $\lim_{r\to\infty}\frac{r^{2}}{32\pi E_{ADM}}\int_{S_{r}(0)}\big(\pi(\nu,\nu)\big)^{2}\,\nu\,d\mu_{\delta}.$
Eichmair and Koerber [willcen] showed that, in general, the center of the Willmore foliation need not converge. However, under more restrictive assumptions on the scalar curvature, the center of the Willmore foliation is well defined and coincides with the Hamiltonian center of mass.
Theorem 3.1 ([willcen, Theorem 2]).
@@
It is then natural to ask whether the center of the foliation by Hawking surfaces is well defined and if it is somehow related to the STCMC center of mass constructed in [STCMC]. To address this, we study the coordinate center of the foliation produced in Theorem 2.10.
The analysis in this section follows the general strategy developed by Eichmair and Koerber [willcen] for the Willmore foliation, adapted here to the setting of area–constrained Hawking spheres. In particular, several steps—such as the quantitative estimates for the center parameter and the control of geometric quantities along the family of surfaces—parallel the corresponding arguments in [willcen], with the necessary modifications to account for the additional kk–terms and the different Euler–Lagrange structure of the Hawking functional. We therefore outline only the main differences and highlight the new estimates specific to the present setting. First we outline the main steps to guide the reader.
$i)$ Improved control of the translation parameter. Under strengthened parity/decay hypotheses we sharpen the bound on the centering parameter $\xi(r)$ so that it cannot drift, this prevents the leaves and their centers from escaping to infinity.
ii)ii) Recentering and removing the constant mode. We rewrite our surfaces as graphs over coordinate spheres with a radius r2r-2. This radius is chosen so that the graph function $u$ has vanishing average; this eliminates the constant mode and isolates the genuine translation information in $\xi(r)$.
$iii)$ Schwarzschild comparison for geometric quantities. geometric data of $\Sigma(r)$ is compared to those of the corresponding coordinate sphere in the Schwarzschild metric. Here most of the $k$ terms can be treated as a perturbation therefore required estimates coincide with the comparison bounds in [willcen] and carry over in our setting.
$iv)$ Expansion for the center and identification of the limit. Finally, again following [willcen] we get a nice expression for $r\xi(r)$ which allow us to study the center of the foliation.
As observed earlier, the foliations need not be centered in general. This indicates that centering requires stronger asymptotic symmetry and decay. Accordingly, we impose the following additional hypotheses on the scalar curvature and the tensor $k$:
@@
These symmetry assumptions are a direct consequence of the Regge-Teitelboim conditions introduced in [regge1974role]. In particular, they guarantee the convergence of the Hamiltonian center of mass. Note also that under (64) the correction term (62) of the STCMC center of mass vanishes.
@@
using $x^{i}\partial_{i}(|x|^{2}R)\leq\mathcal{O}(|x|^{-3})$, the parity conditions and using the same argument as in [willcen, Lemma 11] (or [eichko, Lemma 24]) we have
@@
This estimate is crucial to identify a consistent geometric center for the foliation; without it, the notion of center could drift with the radius.
We have seen that each leaf $\Sigma_{\xi,r}$ of the foliation can be expressed as the graph over $S_{\xi,r}$ of a function $u=-2+\mathcal{O}(|\xi|^{2})$, this implies that when comparing $\Sigma_{\xi,r}$ with $S_{\xi,r}$ the difference would be given by $2+\mathcal{O}(|\xi|^{2})$. The comparison between the two surfaces will be important when trying to study the center of the surfaces $\Sigma_{\xi,r}$ in terms of the center of the spheres $S_{\xi,r}$, that is why it is more practical to consider the spheres $S_{\xi,r}$ with a different radius $\tilde{r}=r-2$ and have the difference between $\Sigma_{\xi,r}$ and $S_{\xi,r}$ be given by the error term $\mathcal{O}(|\xi|^{2})$. Summarizing, following [willcen] we define:
@@
With these definitions and the previous results, we obtain the following result, which is the same as [willcen, Lemma 13] and has the same proof.
@@
We also have the following result equivalent to [willcen, Lemma 14]
@@
Note that (71) is the same expression obtained in [willcen, Lemma 14] ($k$ doesn’t play any role) and (72) is obtained directly by using the decay of $k$. This shows that we can reuse many of the results of [willcen] concerning the expression (71). We can then apply directly [willcen, Proposition 18] to obtain
@@
It was shown in [willcen, Proposition 18] that the second term of the previous expression is related to the Hamiltonian center of mass. To see this, note that the center of mass of a Schwarzschild manifold is given by the zero vector, now since our manifold is asymptotic to Schwarzschild, the center of mass will only depend on the perturbation term of the metric (the tensor $\sigma$). Using the definition of the Hamiltonian center of mas (61) and its linearity we have
$z^{\alpha}:=\frac{1}{32\pi\,r}\int_{S_{r}(0)}\sum_{i,j=1}^{3}\Big(x^{\alpha}x^{j}\big[(\partial_{i}\sigma)(e_{i},e_{j})-(\partial_{j}\sigma)(e_{i},e_{i})\big]-x^{i}\big[\sigma(e_{i},e^{\alpha})-\delta_{i}^{\ \alpha}\,\sigma(e_{i},e_{i})\big]\Big)\,\mu_{\delta},$
for $\alpha=1,2,3$, and $\lim_{r\to\infty}z^{\alpha}=C_{H}^{\alpha}.$ Using integration by parts and the decomposition $e_{\alpha}=e_{\alpha}^{\perp}+e_{\alpha}^{\top}$ with respect to $\delta$, we obtain
(74) $z^{\alpha}=\frac{1}{32\pi\,r}\int_{S_{r}(0)}\big((\partial_{\alpha}\sigma)(\nu,\nu)-\partial_{\alpha}\operatorname{tr}\sigma+\frac{2}{r}\,\operatorname{tr}\sigma\,\nu^{\alpha}\big)\,d\mu_{\delta}.$
Then replacing this second term and the variation $\bar{D}\int_{S_{\tilde{\xi},\tilde{r}}}P^{2}\,d\mu_{\delta}=\int_{S_{\tilde{\xi},\tilde{r}}}2rW_{2}\nu\,d\mu_{\delta}$ into (73), we arrive at the following result which is the dynamical version of [willcen, Corollary 19].
@@
Let $(M,g,k)$ be $\mathcal{C}^{4}$-asymptotic to Schwarzschild with mass $m>0$ satisfying the conditions of Theorem 2.10. Furthermore, assume that
@@
Under the stronger decay and parity hypotheses (and in the time-symmetric case), Theorem 3.1 shows that the center of the Willmore foliation agrees with the Hamiltonian center of mass. In general, the Willmore center is sensitive to the asymptotic distribution of the scalar curvature and may differ from the Hamiltonian center unless additional strong asymptotic symmetry is imposed; this is evident from (75) with $k=0$ and was also observed by Eichmair–Koerber [willcen].
@@
$\begin{split}\frac{r^{4}}{128\pi}\int_{S_{r}(r\xi(r))}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}&=\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}\\ &+\frac{r^{4}}{256\pi}\bar{D}^{2}\left(\int_{S_{r}(\xi r)}2W_{2}\delta(e_{i},\nu)d\mu_{\delta}\right)_{|\xi=\tilde{\xi}}(r\xi-q,r\xi-q)\end{split}$
Where $\tilde{\xi}$ is some intermediate point on the straight‐line segment from $|q|$ to $\xi r$. Now as in Lemma 3.2 we see that the third term decays as $\mathcal{O}(|\xi|)=\mathcal{O}(r^{-1})$ obtaining
$\begin{split}\frac{r^{4}}{128\pi}\int_{S_{r}(r\xi(r))}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}=&\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\end{split}$
@@
(81) $\begin{split}r\delta(\xi(r&),e_{i})=\delta(C_{H},e_{i})+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\mathrm{Sc}\,\delta(e_{i},\nu)\,d\mu_{\delta}\\ &+\frac{r^{3}}{128\pi}\int_{S_{r}(q)}\Big(\delta(e_{i},\nu)\delta(r\xi-q,\nu)\bar{D}_{\nu}\mathrm{Sc}+3\delta(e_{i},\nu)\delta(r\xi-q,\nu)\frac{\mathrm{Sc}}{r}-\frac{\mathrm{Sc}}{r}\,\delta(e_{i},r\xi-q)\Big)d\mu_{\delta}\\ &+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}W_{2}\delta(e_{i},\nu)\,d\mu_{\delta}+\frac{r^{4}}{128\pi}\int_{S_{r}(q)}2\mathcal{D}W_{2}\left(\delta(e_{i},\nu),\delta(r\xi-q,\nu)\right)d\mu_{\delta}+\mathcal{O}(r^{-1})\end{split}$